OpenStudy (anonymous):

hey i have a question, how do you find the implicit differentiation of sin (x/y) = 1/2

OpenStudy (sgvxfcbxvxv):

x=1 and y=2

OpenStudy (anonymous):

let u=x/y ==> sin()u) =1/2 ===> d/du sin(u) = u' cos(u) = (y+y'x)/y^2 cos(x/y)